the perimter of a rectangale is 54cm the length is 3 less than twice the width find the length and the width what is the anwser

Answer by macston(5194)

W=width; Length=L=2W-3; Perimeter=P=2(L+W)
P=2(L+W) Substitute for L
54cm=2((2W-3)+W) Divide each side by 2.
27cm=3W-3 Add 3 to each side.
30cm=3W Divide each side by 3.
10cm=W ANSWER 1: The width is 10cm.
L=2W-3=2(10cm)-3=20cm-3=17cm ANSWER 2: The length is 17 cm.
CHECK:
P=2(L+W)
54cm=2(17cm+10cm)
54cm=2(27cm)
54cm=54cm